### v2

# ViTextVQA: A Large-Scale Visual Question Answering Dataset for Evaluating Vietnamese Text Comprehension in Images

Abstract
Visual Question Answering (VQA) is a complicated task that requires the capability of simultaneously processing natural language and images. Initially, this task was researched, focusing on methods to help machines understand objects and scene contexts in images. However, some text appearing in the image that carries explicit information about the full content of the image is not mentioned. Along with the continuous development of the AI era, there have been many studies on the reading comprehension ability of VQA models in the world. As a developing country, conditions are still limited, and this task is still open in Vietnam. Therefore, we introduce the first large-scale dataset in Vietnamese specializing in the ability to understand text appearing in images, we call it ViTextVQA (Vietnamese Text-based Visual Question Answering dataset) which contains over 16,000 images and over 50,000 questions with answers. Through meticulous experiments with various state-of-the-art models, we uncover the significance of the order in which tokens in OCR text are processed and selected to formulate answers. This finding helped us significantly improve the performance of the baseline models on the ViTextVQA dataset. Our dataset is available at this link for research purposes.
keywords:
Text VQA, Scene Text VQA, OCR Arrangement, Transformer, Vietnamese
1 Introduction
In recent years, Visual Question Answering (VQA) has received special attention from researchers in the fields of Computer Vision (CV) and Natural Language Processing (NLP). The appeal of VQA is boosted when Flamingo [1], GPT4 [2], Gemini [3] - the most powerful chatbots currently integrated with question and answer features on images. Therefore, this field is on a strong development path around the world, especially in rich resource languages such as English and Chinese. Many VQA datasets such as [4, 5, 6, 7] have been released, and new methods with outstanding performance continuously developed, making this multimodal learning with remarkable advancements. Neural and transformer architectures play a pivotal role in a diverse array of tasks that extend beyond traditional methods such as [8, 9, 10].
This task requires a flexible combination of image processing and language processing to address the challenges associated with answering questions based on visual information. In VQA, the model needs to understand not only the content of an image, video, or document image but also the meaning of the question. The model must then generate a meaningful and relevant answer to that question. This places high demands on a deep understanding of both natural language and image analysis capabilities and requires the model to be able to synthesize diverse information from different information sources.
In the context of Vietnamese VQA, Tran et al. [11] built the ViVQA dataset, the first dataset for research on the VQA task for Vietnamese. The ViVQA dataset with 15,000 samples is a small part of the VQA v2 dataset as it was constructed using the semi-automatic method. However, the publication of this dataset is not convincing enough. The shortcomings of the dataset have been elaborately presented by the research team of Nguyen et al. [12]. Therefore, in 2023, they launched a dataset for the open-ended VQA task in Vietnamese called OpenViVQA. This also opens up a new direction of research as well as a different form of the VQA task, which has open-ended questions and open-ended answers in Vietnamese. The dataset includes 11,199 images together with 37,914 question-answer pairs (QAs) created manually.
While the ViVQA dataset does not mention scene text, the OpenViVQA dataset has half of the total samples that exploit this element as one of its challenges. However, OpenViVQA also brings with it some limitations, creating an opportunity for the development of a new dataset that focuses specifically on this aspect.
There are several reasons why OpenViVQA has limitations. First, the number of samples in the dataset may not be large enough to thoroughly and reliably evaluate the performance of the model in scene text. Second, the diversity of scene text in the dataset may be limited, unable to fully reflect the real-life situations the model may encounter. Finally, evaluation capabilities on the OpenViVQA dataset may not reach the level of detail and rigor for scene text processing capabilities when the answers to questions in this dataset are in free form.
Figure 1: Samples extracted from the ViTextVQA dataset.
To address the above limitations, we developed a new dataset called Vietnamese Text-based Visual Question Answering (ViTextVQA), which includes 16,762 images and 50,342 question-answer pairs. This dataset focuses mainly on capturing information from text and scene text appearing in images to address the limitations that traditional VQA models often encounter. ViTextVQA is not only a diverse dataset but also creates a basis for evaluating and improving VQA models capable of handling text, especially in the context of the Vietnamese language.
Our main contributions include the following:
  • Create the first high-quality large-scale dataset for text-based VQA task in Vietnamese, focusing on scene text and text appearing in the image.
  • Analyze the challenge of the ViTextVQA dataset by evaluating the performance of the OCR system.
  • Through our extensive experiments, we found that VQA models using ViT5 as their backbone behave as the answer selector methods when OCR text is suffixed for the question.
  • Demonstrate the important role of OCR text on the performance of VQA models on the ViTextVQA dataset, especially arrangements. Our findings underscore the effectiveness of the top-left to the bottom-right sort, resulting in remarkable enhancements in the performance.
This article is organized explicitly as follows: Section 2 presents an overview of studies related to our research field. In Section 3, we introduce the data creation process and some detailed analysis of our dataset. Next, Section 4 introduces some baselines that we conduct experiments with. Section 5 details the design of the experiments and the evaluation of the results, including comparisons with the baselines. Section 6 analyzes in detail the results achieved. Section 7 discusses some Vietnamese characteristics. Finally, Section 8 summarizes the research and suggests future directions.
2 Related work
2.1 Visual Question Answering datasets
2.1.1 Former Visual Question Answering Datasets
Table 1: Overview of the former VQA datasets in English.
Dataset Type Image Source Annotation Method Language DAQUAR Non text NYU-DepthV2 Human annotation English VQA v1 Non text MS COCO Human annotation English VQA v2 Non text MS COCO Human annotation English TextVQA Text MS COCO Human annotation English OCR-VQA Text Book covers, movie covers Semi-Automation English ST-VQA Text Multiple sources Human annotation English DocVQA Text Scanned documents, receipts etc. Human annotation English VisualMRC Text Internet sources Human annotation English
There are a lot of large-scale and high-quality VQA datasets in the world, especially in the English language. These datasets are not only valuable resources but also crucial sources of inspiration to assist us in constructing the ViTextVQA dataset. Based on research and VQA datasets in English, we can learn and apply effective methods and techniques in the process of developing our Vietnamese VQA dataset. More details of these datasets are in Table 1.
DAQUAR dataset: Malinowski and Fritz [13] introduced the DAtaset for QUestion Answering on Real-world images (DAQUAR) dataset as an early VQA benchmark, initiated the Visual Turing Challenge. They introduced this dataset as a benchmark for evaluating the performance of VQA models. It aimed to assess the capability of models to process both image data and natural language questions, showcasing their ability to bridge the gap between vision and language understanding. The dataset includes a variety of question types, ranging from simple factual queries to more complex questions that necessitate deeper understanding and reasoning about the visual content. This diversity aims to challenge models across different aspects of visual and linguistic comprehension.
VQA v1 dataset: Antol et al. [14] first proposed the free-form question in the VQA task and published the Visual Question Answering v1 (VQA v1) dataset for this task. This dataset has opened a major revolution in natural language processing and has attracted a large number of researchers. During those years, there have been many methods proposed to solve the problem, typically Multimodal Compact Bilinear Pooling (MCB) [15], Multimodal Residual Learning (MUTAN) [16], Neural Module Network (NMN) [17].
VQA v2 dataset: The VQA v1 dataset still has shortcomings that researchers need to pay attention to. The question types were repeated and the answers were only based on simple image features, which led to VQA models at that time only focusing on learning image features and not understanding question content and visual information. To improve visibility for training models, Goyal et al. [18] published the Visual Question Answering v2 (VQA v2) dataset by adding images to the original dataset of Antol et al. [14] and the questions link not just to a single image but to a pair of similar images that lead to two different answers to that question.
TextVQA dataset: Traditional VQA task often ignore text content appearing in images, or they are not capable of reading text appearing in images. To expand the scope of this task, Singh et al. [4] built the TextVQA dataset, which contains 45,336 questions on 28,408 images, requiring questions to use text appearing in the image as the answer. At the same time, they also presented the LoRRA model, which added OCR to retrieve text appearing in images to use as answers.
OCR-VQA dataset: Realize that VQA methods often do not consider text processing in images, which can reduce system performance. To solve this problem, Mishra et al. [19] proposed a new task called Optical Character Recognition - Visual Question Answering (OCR-VQA). The goal of this task is to answer questions about the image by reading and understanding the text it contains. To support research on this task, the article introduces a large dataset called OCR-VQA-200K, including 207,572 book cover images and more than 1 million corresponding question-answer pairs using the semi-automation annotation method.
ST-VQA dataset: VQA datasets and models at the time had serious limitations in ignoring the content of the text in the image. Therefore, when the questions required to understand scene text, VQA models led to disappointing results. To address a significant drawback in VQA datasets and models, [5] introduced the Scene Text VQA (ST-VQA) dataset, specifically designed to tackle the challenge of understanding and answering questions related to texts that appear in the image. The dataset includes question-answer pairs formulated in a way that requires knowledge of the text present in the image for accurate answers. This initiative aimed to improve the performance of models in scenarios where scene text comprehension is crucial.
DocVQA dataset: Unlike existing VQA datasets that focus on natural scene images, DocVQA introduced by Mathew et al. [20] uses images of real-world documents like invoices, receipts, contracts, etc. This presents unique challenges, as the information is often presented in structured formats like tables, forms, and lists, requiring models to reason beyond simple object recognition. DocVQA comprises 50,000 question-answer pairs based on 12,767 document images. These images come from various industry documents and prioritize pages with tables, forms, lists, and figures. This dataset also serves as a benchmark for researchers to develop and evaluate VQA models that can handle document images effectively. Overall, DocVQA provides a valuable resource for advancing research in VQA and its applications in document understanding and information retrieval.
VisualMRC dataset: Tanaka et al. [21] presented a significant advancement in the field of VQA on document images, surpassing the scope of previous datasets like DocVQA by introducing an abstractive novel task. Unlike traditional question-answering datasets, VisualMRC focuses on machine reading comprehension, challenging models to not only extract information from document images but also comprehend and reason about the content to generate abstractive answers. With a substantial dataset comprising over 30,000 question-answer pairs and over 10,000 diverse document images across various web pages, VisualMRC provides a rich and expansive resource for training and evaluating models. The dataset includes open-ended questions, requiring models to generate natural language responses based on a deep understanding of the document layout, text, and visual elements.
2.1.2 Visual Question Answering Datasets in Vietnamese
Table 2: Overview of the former VQA datasets in Vietnamese.
Dataset Type Image Source Annotation Method Language ViVQA Non text MS COCO Semi-Automation Vietnamese EJVVQA Non text Internet sources Human annotation English, Japanese, Vietnamese OpenViVQA Text & Non text Internet sources Human annotation Vietnamese ViTextVQA(ours) Text Internet sources, hand-craft Human annotation Vietnamese
We not only refer to English sources but also to current available VQA Vietnamese datasets when developing ViTextVQA. Referencing several Vietnamese datasets, we better understand the language structure and specific content of Vietnamese, helping to create questions and answers that accurately reflect the cultural context and characteristics of Vietnam. This not only supplements and improves these Vietnamese datasets but also helps overcome their limitations. More details of these datasets are in Table 2.
ViVQA dataset: Tran et al. [11] built the first VQA dataset for the task of automatically asking and answering images based on Vietnamese called Vietnamese Visual Question Answering (ViVQA) dataset. With 10,320 images and 15,000 question-answer pairs, ViVQA meets a large enough number to investigate and develop methods to solve VQA task in Vietnamese. Because the dataset is created using a semi-automatic method, the questions do not really have the nature of the Vietnamese language, and there are no text elements in the image that are mentioned.
EVJVQA dataset: The EVJVQA dataset, introduced in the work by Nguyen et al. [22], stands as a unique compilation of multilingual visual question-answer pairs meticulously curated through human annotation, utilizing images captured in the picturesque landscapes of Vietnam. Encompassing a total of approximately 33,790 question-and-answer pairs alongside a repository of over 5,000 associated images, EVJVQA offers a multifaceted and immersive reservoir of information. Notably, the dataset comprises 11,689 question-answer pairs in Vietnamese, with the remaining entries expertly translated into English and Japanese from this original Vietnamese subset. This expansive dataset has found practical application as the primary data source for the Multilingual Visual Question Answering task in the 9th workshop on Vietnamese Language and Speech Processing (VLSP 2022111https://vlsp.org.vn/vlsp2022/eval/evjvqa), showcasing its significance in advancing research and development in the field.
OpenViVQA dataset: The existing VQA dataset in the Vietnamese language constructed by Tran et al. [11] is addressed as an answer selection or answer classification task. Realize the limitations of ViVQA dataset, Nguyen et al. [12] created the OpenViVQA dataset, the first large-scale dataset for VQA with open-ended answers in Vietnamese, consisting of 11,199 images associated with 37,914 question-answer pairs, In this dataset, answers are created in open-ended answer format, which means the answer is independent of any rule, they are generated in a variety of ways using natural language forms such as words, sentences or phrases. Therefore, this dataset can not be used to conduct experiments as a classification task.
As can be seen in the Vietnamese VQA datasets above reveals certain limitations. Particularly, the ViVQA dataset appears to be merely a translation from English. Furthermore, the EVJVQA dataset does not leverage scene text, thus missing out on potentially valuable information. Evaluating the efficacy of this aspect in the OpenViVQA dataset is challenging due to the nature of its free-form answer. These limitations underscore the further development opportunities, highlighting the potential for improvement in our ViTextVQA dataset.
2.2 Visual Question Answering Methods
In the field of deep learning and the scope of this research, based on VQA models from previous research works, we classify them into three main groups: CNN-RNN, CNN-LM, and ViT-LM. In these groups, CNN-RNN is one of the early and important developments of VQA. CNN-RNN models combine two types of networks: Convolutional Neural Network (CNN) and Recurrent Neural Network (RNN). CNN models are often used to extract features from images, while RNN models are used for natural language processing. When combined, the model is capable of understanding both forms of data and generating or classifying answers to image-based questions. However, using models based on CNN-RNN methods has some limitations, including the inability to effectively handle long questions or questions that are too complex. Therefore, to achieve higher performance and solve increasingly complex challenges, later research has expanded and developed ideas from CNN-RNN. Other branches, such as CNN-LM (combination of CNN and Transformer Language Model) and ViT-LM (combination of Vision Transformer Model and Transformer Language Model), have been proposed to improve the ability to understand and process diverse information from images and questions. However, both ViT-LM-based method and CNN-LM-based method require significant computational resources. For detailed computational resources and training times for models based on these two methods on our ViTextVQA dataset, refer to Section 5. The following Section presents related studies to these types:
2.2.1 Convolutional Neural Network - Recurrent Neural Network Based Method
The deep learning method based on CNN-RNN gained prominence in the years before the advent of transformer architecture. To tackle VQA task, leveraging the strengths of Convolutional Neural Network (CNN) for image understanding and Recurrent Neural Network (RNN) to process sequential information in natural language. In the context of VQA, models that combine CNN with LSTM are employed to extract relevant visual features from the input image and textual features from the input question, respectively. Notable related works include VGGnet [23] and LSTM [24] in VIS+LSTM [25], or another powerful CNN architecture is GoogLeNet [26] and BoW [27] in SMem-VQA [28], or another very deep CNN architecture is ResNet [29] and GRU [30] in Up-Down [31], which have been proven effective in integration facilitates a holistic understanding of both visual and textual modalities, enabling the model to generate accurate answer based on the given image and question.
2.2.2 Convolutional Neural Network - Transformer Language Model Based Method
With the emergence of transformer architecture, especially BERT [32], CNN-RNN based method no longer seems to be the optimal choice. Instead of using RNN, which is limited in capturing the context of the question, Transformer Language Model is becoming a new alternative. CNN-LM based method became attractive for VQA task, combining Convolutional Neural Network (CNN) for image processing and Transformer Language Model to better understand the meaning of the question. Models like ViLBERT [33], VisualBERT [34], Unicoder-VL [35], LXMERT [36], VL-BERT [37], UNITER [38], OSCAR [39] are good examples, they use Faster R-CNN [40] and BERT to embed images and questions respectively, but apply different fusion techniques and attention mechanism to improve model performance.
2.2.3 Vision Transformer Model - Transformer Language Model Based Method
Advances in transformers are not limited to the field of NLP but also open new doors with promising potential in the field of CV. Research by Dosovitskiy et al. [41] demonstrated that the Vision Transformer (ViT) architecture provides impressive results in feature extraction from images, opening up new and promising perspectives for the development of transformers pressure in this field. Models based on ViT and combined with the Transformer Language Model (ViT-LM) emerged as a cutting-edge approach to improve the performance of VQA models. Recent studies have applied this approach and achieved notable achievements, such as BLIP [42], BLIP-2 [43], GIT [44], mPlug [45], Flamingo [1], BeiT-3 [46], PaLI [47] deliver not only outstanding results but also open up new possibilities in practical applications. However, it is essential to note that this advancement often comes at a significant training cost. The learning process of these models requires extensive computational resources and long training time.
3 ViTextVQA Dataset
We conducted a rigorous process of data annotation take over 6 weeks, involving the participation of sixteen native Vietnamese speakers. The process of building the dataset is presented in Section 3.1. The prominent and intriguing features of the dataset, which we have identified through the labeling process, are detailed in Section 3.2. This Section includes an in-depth analysis of elements such as Question, Answer, and Object.
3.1 Dataset Creation
Figure 2: The overview process of creating our dataset ViTextVQA.
The process of creating the ViTextVQA dataset can be outlined as depicted in the diagram presented in Figure 2. After coming up with the idea, we started by collecting images on open platforms such as Google [^2] https://www.google.com/, Facebook [^3] https://www.facebook.com/, Pinterest [^4] https://www.pinterest.com/, etc. and taking photos manually somewhere near the area where we live. Then, pre-processing of the images is applied to remove unwanted images such as duplicates, images without text, etc., and select the images that have the most Vietnamese elements that we can perceive. After that, we divided these images into subsets to facilitate annotation. Once the image preparation process is finished, we then design a special guideline and tool and base it on the annotation process. We also use this guideline for training annotators to start the questions and answers (QAs) creation. The main idea of the guideline is to require the annotators to ask any question related to the image, and the answer must be the text that they see in the image. In the whole process, we hired sixteen ungraduated students for this work, and the Inter-Annotator Agreement was measured twice. This is for evaluating whether the guideline is good or not as well as closely monitoring the annotators to ensure the dataset has the best quality. The next step involves checking and adjusting the dataset, which includes a thorough review of spelling errors and ensuring alignment with the desired quality standards. The final step of this process is to divide all samples into training, development, and test sets.
3.1.1 Image Collection
First of all, we created a list containing many keywords related to images of big cities in Vietnam such as Hanoi, Ho Chi Minh City, Da Nang, Hai Phong, Hoi An, Can Tho, etc. For example, “bảng hiệu ở Thành phố Hồ Chí Minh” (“signs in Ho Chi Minh City”), “khu du lịch ở Hội An” (“tourist areas in Hoi An”), “bảng chỉ dẫn đường ở Hà Nội” (“road signs in Hanoi”, “banner quảng cáo” (“advertising banners”), “bảng tên đường ở Cần Thơ” (“street signs in Can Tho”), and many more another keywords.
To collect image data, we employed Selenium [^5] https://www.selenium.dev/, a widely-used tool for web data crawling, to automatically search and retrieve images from Google Image and Pinterest through the keyword lists we built. We also conducted a manual collection from Facebook and Instagram. However, the performance was not as expected.
To bring more diversity and richness to our image resources, the decision was made to take photos in the neighborhoods where we live. This not only helps us control the quality of our data but also brings great diversity to our datasets.
By the end, we collected more than 40,000 bearing beauty and characteristic images in Vietnam. However, crawling this data cannot avoid noise and unwanted images. For duplicate images, we use the hash algorithm to remove them, but we still cannot completely resolve them because some images have different resolutions but are same image. Therefore, we encourage annotators to make a strong decision to remove images that they feel they have previously annotated. To filter images containing scene text, we used SwinTextSpotter [48] - one of the powerful OCR tools that supports Vietnamese.
Finally, we obtained more than 22,000 high-quality images. To facilitate the annotation process, we organized these images into many different subset folders, with each folder containing about 300 images.
3.1.2 Annotation Tool
To increase the convenience and efficiency of data annotation, we have developed a labeling support tool specifically for the ViTextVQA dataset named “VQALabeling” [^6] https://github.com/hieunghia-pat/VQALabeling.git. Utilizing the Qt framework and C++ programming language, we crafted a desktop application that operates seamlessly across multiple platforms including Windows, Ubuntu, and CentOS. This tool provides a user-friendly interface (see Figure 4), which is designed to facilitate efficient data management at the image folder level, ensuring organized and effective annotation for the ViTextVQA dataset. Upon completion of annotating a folder of images, the associated annotations information is stored in a *.json file (see Figure 4) structured as follows:
  • annotations: List of question-answer pairs of images. Each question-answer pair includes the corresponding question and answer.
  • delete: During the image collection process, some images, although meeting size requirements, are not suitable due to being broken, blurred, or cannot be used to ask questions, etc. We use this field to mark whether an image is used in the official dataset or not.
  • filename: Name of the image.
  • filepath: The folder path containing the images.
Figure 3: User interface of VQALabeling.
Figure 4: *.json file storage structure.
3.1.3 Question-Answer Creation
After completing the image collection process, in the early stages of the annotation process, we developed a guideline for annotating. The goal of the guideline is to ensure the validity of question-answer pairs annotated with images. This is a difficult stage, requiring focus and effort because the guideline must be continuously updated throughout the annotation process. We pay special attention to handling unexpected situations and perform Inter-Annotator Agreement measurements periodically to ensure dataset quality. Furthermore, we utilize the “VQALabeling” outlined in Section 3.1.2 for creating question-answer pairs on different images.
General Rules
General Rules are established to ensure that the process of creating questions and answers for images is done carefully and systematically. This regulation helps ensure consistency and quality of data, from ensuring the number of questions and answers per image.
  • Rule 1: For each subset, the annotator created at least 600 pairs of questions and answers.
  • Rule 2: The annotator is encouraged to use their personal vocabulary to ask questions.
  • Rule 3: Each image contains at most 10 pairs of questions and answers.
  • Rule 4: All letters are in lowercase and in the scope of the standard keyboard.
Rules for creating questions
The rules for creating questions provide specific instructions on how to ask questions so that they are specific, not vague, and ensure that the question leads to an answer that is already in the image. This way, not only does it create valuable questions, but it also creates convenience for participants to think and reason creatively and logically.
  • Rule 1: Question must be answered by text that appears only in the image.
  • Rule 2: Do not ask questions with mathematical implications.
  • Rule 3: Do not ask where the text is located.
  • Rule 4: Do not ask options (yes/no).
  • Rule 5: Do not ask the color of the text.
  • Rule 6: Do not ask the number of text that the annotator has to count.
  • Rule 7: Avoid asking questions that are too short.
  • Rule 8: Avoid asking multiple questions that have the same answer in the same image.
  • Rule 9: Avoid asking general questions that lead to many different answers.
  • Rule 10: Take the direction of the annotator seeing the image as the reference.
Rules for creating answers
The rule for creating answers guides annotators to make sure that annotated answers are correct and do not contain extraneous information. This ensures the consistency of the dataset and makes the models easier to process.
  • Rule 1: The answer is only the words that appear in the image.
  • Rule 2: Keep the content intact, convert everything to lowercase.
  • Rule 3: If the text has a line, replace it with a space.
  • Rule 4: If the answer has some words that are obscured, only annotate what the annotator sees.
Next, we recruited sixteen undergraduate students. Then, we trained them to understand the annotation guideline and asked them to annotate on about 100 test images. After ensuring that all students understood and applied the annotation guideline correctly, we appointed one person from our team for every four students to act as a supervisor. Annotation supervisors are responsible for monitoring and supporting the annotators, ensuring that everyone is performing according to the required process and quality.
During the annotation process, we also measured the Inter-Annotator Agreement which is mentioned in Section 3.1.4. If they meet the requirements, the data they annotated will be checked by annotation supervisors for spelling errors and standardized with the same font, punctuation, etc. If not, we will retrain them to follow the guideline to avoid unwanted errors and show them samples that need fixing.
Finally, after six weeks of non-stop work, we and our passionate annotators completed more than 50,000 samples. Then, all annotated data were randomly split into training, development, and test sets corresponding to the ratio of 7:1:2 for future evaluation and analysis.
3.1.4 Inter-Annotator Agreement
When measuring the Inter-Annotator Agreement, our team, those who understand the guideline best, carefully annotate 100 question-answer pairs on about 30 images. Next, we hide the answers in these samples and ask annotators to re-annotate these hidden answers. This was done regularly every two weeks during the annotation process. Due to the nature of this VQA task, we are not aiming for a cheat answer like a multi-label classification task, so the dataset does not have a specific label space. That makes it impossible to measure the Inter-Annotator Agreement using traditional methods such as Cohen’s kappa [49], Fleiss’s kappa [50], and F1-score [51]. Therefore, to address this problem, we choose the F1-score commonly used in the Question Answering task as [52].
In this case, F1-score is computed over the individual words in the prediction(answer annotated by annotators) against those in the true answer (answer annotated by annotation supervisors). The number of shared words between the prediction and the truth is the basis of this metric. F1-score includes precision is the ratio of the number of shared words to the total number of words in the prediction, and recall is the ratio of the number of shared words to the total number of words in the ground truth. The F1-score formula is defined in Section 4.5.2. The requirement we set for annotators was an F1 score greater than 70%.
Table 3: Results of two rounds of Inter-Annotator Agreement measurements.
First Time Second Time
min F1-score (%) 73.91 83.36
max F1-score (%) 87.73 87.05
avg F1-score (%) 81.83 85.34
After measuring the Inter-Annotator Agreement through the F1-score twice, we simplify the results into Table 3. Based on the table, it can be seen that in both evaluations, the lowest F1-score value is relatively high (73.91 and 83.36), this shows that the level of uniformity between annotators is not low. In addition, the average F1-score also increased from the first to the second time, from 81.83 to 85.34, showing an improvement in consistency between annotators. This suggests that the annotated dataset has improved in quality over sessions, and this measurement can be considered reliable to confirm the quality of the dataset.
3.2 Dataset Analysis
Figure 5: Statistic of images and question-answer pairs.
The ViTextVQA dataset includes 16,762 images accompanied by 50,342 QA pairs. Figure 5 depicts the size of training, development and test sets in our dataset. This dataset offers a diverse range of visual scenes as well as scene texts and corresponding questions and answers. Each image is annotated with some question-answer pairs, providing valuable insights into scene text and object required for accurate model predictions where the answer cannot be outside OCR text in the image and this makes the dataset more special in the VQA task. Moreover, we conducted statistics to compare our dataset with other common datasets in this field, and more details can be seen in Table 4.
Table 4: Comparison between well-known VQA datasets.
Dataset Language Images Questions Answers
TextVQA [4] 28,408 45,336 453,360
ST-VQA [5] 23,038 31,791 31,791
DocVQA [20] 12,767 50, 000 50, 000
OCR-VQA-200k [19] 207,572 1,002,146 1,002,146
InfographicVQA [53] 5, 485 30,035 30,035
VisualMRC [21] 10,197 30,562 30,562
VizWiz [54] 32,842 265,420 265,420
OK-VQA [55] 14,031 14,055 14,055
Visual Genome [56] 108,077 1,773,258 1,773,258
UIT-EVJVQA [22] Multilingual 4,879 33,790 33,790
MCVQA [57] - 369,861 369,861
ViVQA [11] Vietnamese 10,328 15,000 15,000
OpenViVQA [12] 11,199 21,271 21,271
ViCLEVR [58] 26,000 30,000 30,000
ViTextVQA (Ours) 16,762 50,342 50,342
To further analyze the characteristics of the ViTextVQA dataset, we compiled statistics on the distribution of the lengths of questions and answers. In addition to that, we also gathered information on Part-of-Speech (POS) and Named Entity Recognition (NER). POS analysis reveals the grammatical structure and meaning of both questions and answers, while NER provides an overview of how words are used in questions and answers. Therefore, we utilized VnCoreNLP [59] - a powerful tool for processing Vietnamese text - to extract POS tags and NER tags. Our statistical overview is presented in Section 3.2.1 and Section 3.2.2.
3.2.1 Question Analysis
Question Length
We performed a statistical examination of the question lengths in Figure 6. In the dataset, question lengths are not only concentrated in a fixed range but are distributed fairly wide range, from very short of only 3 tokens to very long of up to 31 tokens. This creates a considerable variety in terms of question length. The short questions are used often which aim to ask about basic information. On the contrary, there are also questions with considerable length, up to 31 tokens. These questions often involve complex issues, requiring a level of detailed information or multiple interrelated issues. On average, the question length was 9.59 tokens indicating that most questions were not too short but not too long either. This may reflect the attempt of annotators to provide enough information without making the question too cumbersome.
Figure 6: Distribution of question length.
POS Tags in Questions
In VQA tasks, analyzing the Part-of-Speech (POS) tags of words in questions is crucial for understanding their grammatical structure and meaning, providing insights into their overall grammatical makeup.
Figure 7: Distribution of Postag in question.
Figure 8: Distribution of NER in question.
As can be seen in Figure 8, our dataset is annotated using various types of POS tags in question. Among them, Noun (N) appears the most, with more than 140,000 cases, usually referring to objects or people in the image. The Pronoun (P) is used to replace the noun and appears over 61,000 times, playing an important role in expressing the relationship between the components in the question.
Verb (V) is another important part of questions, with approximately 75,000 cases, often describing the action or state of the object in the image. Additionally, the Preposition (E) appears less than half that of the Verb (V) approximately 35,000 times, commonly used to determine relationships between location, time, etc. between the components in the question.
Adjectives (A) and Classifier noun (Nc) appear relatively frequently, with both types of words appearing more than 13,000 times. In the Vietnamese language, Nc represents nouns with classifier properties, while Adjectives (A) are used as a description of their characteristics. The Adverb (R) appears just more than 6,100 times, often used to add information about the manner, extent, time, or frequency of an action, event, or adjective to make the question more vivid.
Other types of words such as Numeral/Quantity (M), Determiner (L), Subordinating conjunction (C), etc. also play an important role in identifying and describing details of images. Although they only make up a small part of the text, they help annotators create diverse grammatical structures and create challenges for VQA models.
NER Tags in Question
Like POS tags, Named Entity Recognition (NER) also provides an overview of language usage in questions, with special attention to entities such as places, organizations, personal entities, and other entities.
Figure 8 shows that the majority of the words in the question of the dataset are not classified as specific entities with more than 350,000 “O” tags which may be common words related to objects such as “person”, “object”, or “action”. Additionally, specific entities such as locations (B-LOC), organizations (B-ORG, I-ORG), individuals (B-PER), and other entities (B-MISC, I-MISC) account for a very small part. This difference can be explained by the fact that questions are not too long and often ask about only part of the text appearing in the image.
3.2.2 Answer Analysis
Answer Length
Figure 9: Distribution of answer length.
Figure 9 shows a large variation in the length of the answers in the dataset, from very short of only 1 token to very long of up to 52 tokens. This creates a significant difference in length between the answers. The average length of each answer is approximately 4.18 tokens, showing that the majority of answers are short. The reason for this phenomenon can be explained by the fact that annotators are constrained not to use tokens that do not appear in the image, leading to answers focusing mainly on from 1 to 3 tokens.
POS Tags in Answer
Similar to the question, POS tags in the answer not only help to better understand the grammatical structure but also provide important information about the content and meaning of the answer in the VQA task.
Figure 10: Distribution of Postag in answer.
Figure 11: Distribution of NER in answer.
In Figure 11, we can notice the richness and diversity of the Vietnamese language used in the answer. Noun (N) make up the largest proportion of the data with approximately 70,000 instances, these tags often appear to designate objects, people, or objects associated with the image. This shows a focus on providing information about what answers the image-related question.
In addition, the Verb (V) also plays an important role. With nearly 30,000 and ranked after noun (N), it shows that annotators are also interested in describing the actions or states of objects in the image, making the answers more vivid.
Punctuation and symbols (CH) and Numeral/Quantity (M) appear the third most with about 25,000 times. This is understandable because the dataset focuses on text in the image, and dots often appear in phone numbers (e.g., 0909.888.888), building addresses with punctuation, or on advertising banners. In contrast, answers about the price of product, number of houses, etc. often contain a numeral or quantity factor (M).
Adjectives (A) are often used to express the characteristics or conditions of the object or entity described in the answer. With around 9,000 occurrences, they show that, even if they are just tokens from the OCR that the annotator sees in the image, they still contain specific messages and linguistic colors to convey meaning best.
Moreover, several other tags, while they may be in the minority, significantly contribute to enriching the diversity of answers in the dataset.
NER Tags in Answer
NER tags also play an important role in identifying specific entities mentioned in the answer, help us better understand the variety of entities emphasized in answering image-based questions, and provide information that the model can pay attention to.
Figure 11 provides insight into how language is used in answers annotated by native Vietnamese speakers, with particular attention to entities such as places, organizations, and individuals. The majority of words in the answer are not classified as concrete entities (O) similar to the question. However, we also see the emergence of specific entities such as places (B-LOC, I-LOC), organizations (B-ORG, I-ORG), and individuals (B-PER), although not too prominent with less than 4,000 or even less. This diversifies the content of the answer, helping the model understand and generalize about possible real-life human cases.
3.2.3 Object Analysis
Object in Image
By Utilizing VinVL [60], we extracted various details about objects within the image, including location coordinates, object names, attributes, and more. This object information enhances our understanding of the visual content depicted in the image.
Figure 12: Distribution of the 50 most frequent objects in images.
Figure 12 shows that “person” ranks first with an occurrence count of over 47,000, showing a trend that most images in our dataset contain human presence. This deeply reflects the importance of humans in the context of the dataset. It can be said that human play a central and important role in the images we collect.
Additionally, “sign” and “letter” ranked second and third with about 43,000 and 36,000 appearances, respectively. This highlights the importance of signs and letters in the dataset, emphasizing our focus on scene text.
Our dataset not only reflects the diversity of people but also the diversity of landscapes, with objects such as “building”, “tree”, “sky”, “street”, and “sidewalk” appeared popular. Each of these objects has thousands of occurrences, creating a rich and diverse scene and highlighting deep and multidimensional images in the dataset.
Besides, “shirt”, “trousers”, “hat”, and “jacket” are objects representing clothing and accessories that appear quite frequently, demonstrating the importance of this element in the dataset. This could be related to fashion, personal style, or the specific setting in the images.
In our dataset, a particularly striking and interesting point about the culture and characteristics of Vietnamese people is the popularity of motorbikes compared to cars. Specific Figure 12 shows that the presence of “motorbike” is significantly more than “car”. With nearly 9,000 occurrences, motorbikes far outnumber more than 3,000 occurrences of cars. This highlights a notable ratio, with motorbikes constituting approximately three times the frequency of cars in the dataset and it suggests that the priority and preference for using motorbikes in daily travel and in the transportation system as well as partly reflecting the development of current Vietnamese culture and lifestyle.
Object in Question
We employed VNCoreNLP [59] to extract POS tags from questions, enabling us to compile statistics on the frequency of mentioned objects. This facilitates a deeper analysis of the questions, shedding light on the prevalence and significance of specific objects in questions of the dataset.
Figure 13: Distribution of the 50 most frequent objects in questions.
The detailed data in Figure 13 clearly illustrates the variety and richness of information that the annotator is interested in “store” and “name” appear especially frequently, more than 8,000 and nearly 7,000 times, respectively, highlighting their interest in specifically identifying the store and the names of the objects in the image.
Objects related to visual content such as “photo”, and “image” appear quite a lot, with a total of about 10,000 occurrences for both, showing special interest in with details of the image. At the same time, attention to information about “number” and “price” is also clearly shown through more than 4,500 and 1,500 appearances, respectively.
Culinary object group such as “eatery” and “dish” attract interest, especially with the desire to learn about the cuisine in the photo. At the same time, group of objects about people such as “person”, “man”, and “woman” also play an important role with a relatively high number of appearances.
Meanwhile, in addition to “store”, information about shopping and trading locations is shown through the significant appearance of “shop”, “supermarket” and “market”, highlighting the interest of the annotators with places in the image.
4 Methodology
Figure 14: The overview of text-based VQA models structure.
The common text-based VQA approach employs three key components to tackle image-based questions: Visual Features Extraction, Textual Features Extraction, and OCR System (see Figure 14). In this framework, the Visual Features Extraction extracts vital information from images, while the Textual Features Extraction encapsulates questions into features vectors. Meanwhile, OCR System undertakes the crucial task of recognizing and extracting text portrayed in images. Through the seamless integration of these three components, the text-based VQA method can comprehend and answer questions from images with enhanced accuracy and efficiency.
4.1 Textual Features Extraction
To achieve a better match with the Vietnamese language and simultaneously enhance performance on the dataset, we opted for pre-trained language models to extract textual features on questions and OCR text for baseline models. These pre-trained models not only capture the nuances of the language but also enhance the capacity to process and comprehend text deeply within the Vietnamese context. Refer the Table 5 mentioned in Section 4.4 to see specific pre-trained language models utilized by baseline models. Pre-trained language models that we chose as follows:
ViT5: The Vietnamese Text-to-Text Transformer named ViT5 in the paper [61], stands as a cutting-edge, preeminent model in the realm of Transformer-based encoder-decoder architectures tailored specifically for the Vietnamese language. Conceived and crafted by VietAI, this state-of-the-art model undergoes extensive training on an extensive and varied corpus of high-quality Vietnamese texts, employing a T5-style self-supervised pretraining methodology. This allows ViT5 to generate text, translate languages, write different kinds of creative content, and answer your questions in an informative way.
mBert-cased: BERT multilingual base cased [62] is a variant of BERT [32], this is a pre-trained multilingual Transformer-based masked language modeling (MLM) model developed by Google. It is a cased model, which means that it distinguishes between different cases of letters, such as uppercase and lowercase. The model is pre-trained on a massive corpus of text data, including Wikipedia articles, books, and news articles in 104 different languages, including Vietnamese. This training allows the model to capture the nuances of multiple languages and perform well on a variety of downstream tasks in multiple languages.
4.2 Visual Features Extraction
ViT introduced by Dosovitskiy et al. [41], a pre-trained image model based on transformer architecture, has been trained on a large amount of image data, playing an important role in the English baseline models mentioned in Section 4.4. In the baseline adaptation to Vietnamese, we kept the same visual feature extraction using ViT. Meanwhile, When adapting Vietnamese to conduct experiments on the ViTextVQA dataset with baseline models based on CNN-LM based method, we chose VinVL [60], a model based on Faster R-CNN [40] to extract object features. This choice also brings convenience when we use VinVL to analyze object in Section 3.2.3.
ViT: Short for Vision Transformer [41] is a new approach to image recognition that utilizes the Transformer architecture. The first step involves dividing the image into smaller squares called patches. These patches are fed into a Transformer encoder, similar to those used in translating languages. This excels at representing relationships between different parts of the input image and helps the model understand the spatial context between various objects. Finally, the rich features are used to classify the image, determining what the image contains. ViT has shown promising results in image processing tasks, achieving superior performance to traditional CNN while potentially requiring less training data.
VinVL: Prior research mainly concentrated on perfecting the way these models fuse visual and language information, neglecting advancements in object detection itself. The authors proposed VinVL [60], a system that emphasizes better visual representations by enhancing the object detection model. They develop a new object detection model with a broader range of object and attribute categories and training on a larger dataset. This improved object detection model is trained on a significantly larger dataset, allowing it to capture more nuanced visual details. VinVL achieves state-of-the-art performance on various VL tasks, including Visual Question Answering (VQA), Image Captioning, Visual Commonsense Reasoning (VCR), etc. VinVL highlights the importance of strong visual representations in Vision Language models. By prioritizing advancements in object detection, the model achieved superior performance across different vision language task.
4.3 OCR system
In contrast to commercial closed source English OCR systems for VQA baseline models which are introduced in Section 4.4 such as Rosetta-en, Amazon OCR, etc., SwinTextSpotter [48] stands out for its open source Vietnamese language support. This capability makes it an ideal tool for extracting vital OCR features tailored specifically to the Vietnamese language. Consequently, we opted for SwinTextSpotter to extract essential OCR features for our baseline models adapted for Vietnamese. This choice ensures that our baseline models are well-equipped for better performance in Vietnamese text.
SwinTextSpotter: Current state-of-the-art OCR approaches typically share a common backbone for both text detection and recognition tasks. However, this does not fully exploit the potential interaction between the two. Therefore, the authors introduced a new framework named SwinTextSpotter [48]. It is an end-to-end system that aims to achieve better synergy between detection and recognition. SwinTextSpotter achieves better performance in scene text spotting tasks compared to previous methods in many different languages, including Vietnamese.
4.4 Baseline Models
Figure 15: Timeline of Visual Question Answering method.
There are numerous effective methods have emerged globally for addressing VQA task. To evaluate the difficulty of the ViTextVQA dataset, we choose M4C [63], LaTr [8], PreSTU [10], BLIP-2 [43] and SaL [64] according to the historical development (Figure 15) of VQA methods and use them as baselines to perform experiments. Note that these baselines were originally designed to support English, but we adapted them to Vietnamese while still retaining their essence.
Table 5: Features extraction and OCR system of baseline models.
Model Visual Features Textual Features OCR System Method Type
M4C VinVL mBERT SwinTextSpotter CNN-LM
BLIP-2 ViT ViT5 SwinTextSpotter ViT-LM
LaTr ViT ViT5 SwinTextSpotter ViT-LM
SaL VinVL ViT5 SwinTextSpotter CNN-LM
PreSTU ViT ViT5 SwinTextSpotter ViT-LM
For a comprehensive overview of the baseline models, refer to Table 5 along with additional details provided in Sections 4.1 and 4.2 regarding their pre-trained model. Furthermore, Section 4.3 details the OCR system, and Section 2.2 elaborates on the different types of methods employed by these models.
M4C: The emergence of Facebook’s TextVQA dataset [4] catalyzed the development of a pioneering approach called Multimodal Multi-Copy Mesh (M4C) [63], marking the first instance of treating VQA as a generated answer. M4C employs BERT [32] for question embedding, utilizing its encoder module as a multimodal encoder to embed diverse features, including object and region features in images, question features from BERT, visual features of scene texts, and features of previous tokens in answers. In a departure from conventional classification approaches, M4C views VQA as a generative task and introduces the Dynamic Pointer Network, as proposed by Hu et al. [63]. This network seamlessly integrates features of embedded tokens in the vocabulary with those of scene texts. In particular, M4C makes decisions at each step of the answer generation process, determining whether to select tokens from the vocabulary or scene texts.
LaTr: The Layout-Aware Transformer (LaTr), as introduced by Biten et al. [8], adopts the encoder-decoder transformer architecture, specifically building upon T5 [65]. The model comprises three primary blocks. The initial module is dedicated to a Language Model trained for document layouts, focusing on text and layout details. The second module employs spatial embedding to embed tokens related to scene text along with their positional information, represented by OCR tokens bounding boxes. Utilizing the Vision Transformer (ViT) [41], the third module extracts visual features. Subsequently, these embedding features from the three modalities serve as input for the pre-trained transformer. The encoder then learns a representation that harmonizes information from these modalities. This learned representation is subsequently employed by a decoder for analysis and generation of an output, typically in the form of an answer.
PreSTU: Pre-Training for Scene-Text Understanding [10] stands out from other models in tackling a common task by incorporating a unique approach. To ensure that models grasp spatial information in scene text and standardize the target output sequence during training, the authors arranged OCR text by their positions, sequentially from the top-left to the bottom-right. They then concatenated the sorted texts using a T5 separator ¡/s¿. Afterward, they randomly split the OCR text into two segments, utilizing the first part as additional input and the second part as the target. Next, the research team applied the T5 architecture to conduct pre-training on a large amount of scene text image data. The goal is to help the model, which is a language model, be able to learn information related to scene text in the image. During the pretraining and finetuning phase, ViT [41] is used to extract image features, assisting PreSTU in predicting the target answer based on the question and OCR text contained in the image.
BLIP-2: BLIP-2 [43] short for “Bootstrapping Language-Image Pre-training with Frozen Image Encoders and Large Language Models”, is a model that can understand and generate text related to images and do so in languages. It combines a vision transformer, a query transformer, and a large language model (LLM). In the structure of the model, they frozen image encoders and large language models (LLMs), meaning these components do not learn during training, query transformer (Q-Former) that acts as a lightweight bridge between the image encoder and LLM. This model was trained using the weights of ViT and T5 on multiple VQA datasets.
SaL: Most current Text-based VQA methods focus on improving network design or pre-training tasks. However, these approaches often treat text in images as a simple sequence, neglecting the lack of semantic connections between many words and the complex spatial relationships between them. Fang et al. [64] proposed a novel method called “Separate and Locate” (SaL) that addresses these issues. SaL explores two special modules, including “Text Semantic Separation” (TSS) and “Spatial Circle Position” (SCP). TSS module identifies words that have true semantic relationships and separates them from those that are simply positioned close together. This helps the model focus on meaningful connections between words. SCP module introduced a new way to represent the spatial relationships between words. Instead of relying on simple 1D positioning, SCP uses a “spatial circle” embedding that captures the relative distances and directions between words, providing a more accurate understanding of their layout. By combining these two modules, SaL achieves significant improvements in accuracy on TextVQA [4] and ST-VQA [5] datasets, outperforming state-of-the-art methods without requiring any pre-training itself.
4.5 Evaluation Metrics
4.5.1 Exact Match
Exact Match (EM) score is a metric commonly used in NLP tasks, particularly in Question Answering systems. It measures the percentage of predictions made by a model that exactly match the target or ground truth answers.
In the context of question answering, an answer is considered to be correct only if it matches the reference answer exactly. The EM score is a straightforward evaluation metric that provides a clear indication of how often a model produces answers that are exactly right.
The formula for EM score is:
EM=Count of questions predicted exactlyCount of $\textbf{EM}=\frac{\text{Count of questions predicted exactly}}{\text{Count of questions}}$ (1)
4.5.2 F1-score
F1-score is a common way to measure how well a system performs in question-answering task. It looks at how much the answers of system match the correct ones. Precision focuses on the accuracy of the answers of system, Recall looks at how many correct answers the system can find, and F1 combines both precision and recall for a more comprehensive evaluation. The purpose of this metric is to see how good a system is at providing accurate and complete answers.
The formula for Precision, Recall, and F1 Score are determined as follows:
Precision=Count of tokens predicted exactlyCount of tokens $\textbf{Precision}=\frac{\text{Count of tokens predicted exactly}}{\text{Countof tokens predicted}}$ (2)
Recall=Count of tokens predicted exactlyCount of gold standard $\textbf{Recall}=\frac{\text{Count of tokens predicted exactly}}{\text{Count ofgold standard tokens}}$ (3)
[…]
5 Experiment and Results
5.1 Experimental Configuration
The baseline models were trained using the Adam optimization algorithm [66] on an RTX 4090 GPU with 24GB of memory. Each model underwent 10 epochs, with approximately 7 hours required for completion using ViT-LM based method and 6 hours using CNN-LM based method. Hyperparameters for the baseline model were set as follows: a learning rate of 3.0e-5, dropout rate of 0.2, batch size of 16, and weight decay of 1.0e-4.
For baseline models with large version, training was extended to 10 hours over 10 epochs using ViT-LM based method and 9 hours with CNN-LM based method. All other hyperparameters remained consistent with the configuration in base version.
5.2 Main Results
Table 6: Results of models on ViTextVQA test set.
Model Version Method Type F1-Score (%) EM (%)
M4C base CNN-LM 25.46 10.26
BLIP-2 base ViT-LM 36.87 13.59
BLIP-2 large ViT-LM 38.47 15.58
LaTr base ViT-LM 40.41 17.85
LaTr large ViT-LM 42.34 20.94
SaL base CNN-LM 43.49 20.11
SaL large CNN-LM 43.39 21.48
PreSTU base ViT-LM 43.81 20.85
PreSTU large ViT-LM 44.93 22.64
Table 6 presents a comprehensive of the results achieved by individual models, showcasing their F1-Scores and EM metrics. The findings underscore significant disparities in performance among the various models, providing valuable insights into their respective strengths and weaknesses. Notably, within the baseline models, PreSTU large version based on ViT-LM method stands out as the top performer, showcasing an impressive F1-Score of 44.93 and an EM of 22.64. Moreover, the table illuminates the substantial performance enhancements attainable in VQA task through the utilization of large-scale pre-trained models.
6 Results Analysis
6.1 ViT5 and Effect of OCR Text as Question Suffix
During the experiment, we discovered that models using ViT5 as their backbone achieve better performance when the given question is suffixed with OCR text from images compared to the scenario in which only the question is passed to the VQA models as in previous studies. In this respect, we are interested in further exploring the performance that models such as BLIP-2, LaTR, PreSTU, and SaL perform when using OCR text as an important part of the “context” for the question. This experiment was conducted by adding and removing OCR text along with the question, thereby helping us with more detailed assessment of the capabilities of the models.
Figure 16: Performance of models include OCR text and without OCR text scenario.
Figure 16 (left) illustrates the F1-score of the models when using and not using OCR text. Models trained with question extended by OCR text consistently outperformed models that did not use OCR text. In particular, PreSTU model (large) is a standout, with an F1 score increasing from 16.95 (without OCR) to 44.93 (with OCR), showing a significant increase in performance when integrating OCR text in the training process. This positive trend is also repeated in other models.
Similarly, with the EM score which is shown in Figure 16 (right), it can be seen that models trained with OCR text consistently demonstrated almost twice the accuracy compared to models that did not use OCR. This indicates that integrating OCR text into the training process not only improves performance but also enhances the accuracy of the models.
Overall, the use of OCR text to expand the question as a “context” for models using ViT5 as their backbone has proven effective and has the potential to improve the performance of these models on VQA task. In other words, as answers in the ViTextVQA dataset are text or scene text from images, providing OCR text along with questions makes the VQA models in our experiments find it easier to “select” appropriate tokens from OCR text for determining the answers.
6.2 Effect of OCR System Performance and Challenge
Figure 17: The ratio of percentage answer token appears in the OCR text in test set.
Based on experiments, we observed a correlation between the performance of the OCR system SwinTextSpotter and the efficacy of the models. To analyze this relationship, we partitioned the test set data based on the percentage of answer tokens present in the OCR text extracted by SwinTextSpotter. The evaluation was carried out on segments with varying ratios: less than 25%, 25% to less than 50%, 50% to less than 75%, 75% to less than 100%, and 100%. Refer to Figure 17 for a detailed breakdown of these ratios.
To evaluate the impact of the performance of the OCR system on the overall performance of the VQA models, we conduct a detailed analysis in the context of both the VQA model and the ratio appearance of answer tokens in OCR text extracted by SwinTextSpotter.
Figure 18 clearly depicts that despite the diversity in the models, the general trend in F1-score is similar. Performance increases as the number of answer tokens appearing in the OCR text increases. This clearly demonstrates the importance of information in OCR text, providing important context and supporting data for the VQA model to predict accurately.
Figure 18: Results of models based on OCR system performance.
Figure 18 also illustrates in detail the EM results of the basic models on the occurrence rates of answer tokens in OCR text. It shows a general zig-zag trend, where performance drops when the proportion of answer tokens appearing in OCR text is between 25% and below 50%. Performance increases as the token ratio increases from 50% to less than 75%, but decreases when the token ratio is 75% to less than 100%. In particular, performance increases sharply when the ratio reaches 100%.
It should be noted that, even with a 100% rate of answer tokens appearing in OCR text, the EM and F1-score performances are still only low, lower 35.00 and lower 60.00 respectively. This highlights the major challenge that VQA models face with the ViTextVQA dataset, which requires the ability to synthesize information from multiple sources and the ability to remove noise from OCR text to be able to predict answer correctly.
6.3 Effect of Answer and Question Length
Table 7: Group of answer and question length in test set.
Group Length (n) Answer Samples Question Samples
Short $n\leq 5$ 7853 225
Medium $5<n\leq 10$ 1541 6834
Long $10<n\leq 15$ 421 2617
Very long $n>15$ 213 352
We divided questions and answers based on their length in the number of tokens. This classification helps us evaluate the performance of the underlying models more flexibly and accurately. The number of samples in the test set based on different lengths is detailed in Table 7. Classification is done as follows:
  • Short question (and short answer): These are questions and answers that do not exceed 5 tokens in length. These are usually simple questions with concise answers.
  • Medium question (and medium answer): This group includes questions and answers ranging from 6 to 10 tokens in length. Questions can be more complex, requiring more detailed information, while answers provide complex answers concisely.
  • Long question (and long answer): Questions and answers in this group range from 11 to 15 tokens in length. These questions require detailed information or are more complex. Answers in this group also contain rich and detailed information.
  • Very long question (and very long answer): The last group includes questions and answers that are more than 15 tokens. These are often very complex questions, requiring models with strong generality and performance to handle them. Answers in this group also contain a lot of detailed information about the context and details in the image.
Figure 19: The results of models based on answer length.
Figure 19 illustrates that models have the same trend in performance with different answer lengths, from short to very long. The results from models clearly demonstrate that as the length of the answer increases, there is a corresponding decrease in performance, as evidenced by a gradual reduction in EM scores. Conversely, with the F1 metric, short answers do not achieve the highest scores like EM. For medium and long answers, the F1 scores tend to be higher. However, this does not mean that longer answers always result in higher F1-score. It clearly shows that the longer the answer, the more F1-score decreases.
Figure 20: The results of models based on question length.
As can be seen in Figure 20, in terms of the F1-score metric, the highest scores are observed for short questions, gradually decreasing as the question length increases and reaching the lowest scores for very long questions across all models. However, for the EM metric, the performance is relatively stable across different question lengths, from short to long. Notably, the SAL base and PreSTU base exhibit different behaviors. While the SAL base appears to increase in performance as the question length increases from short to long, the PreSTU base shows a significant decrease. Particularly, the performance of models becomes increasingly challenging as the question length becomes very long, resulting in a notable decrease in scores.
6.4 Effect of OCR text size
Figure 21: Samples of OCR text size group.
We evaluated the performance of the models on different OCR text sizes to better understand this aspect. To ensure the accuracy of the evaluation, we need to select samples where all tokens in the answer appear in the OCR text. This is important because if any token falls outside the range of recognizable OCR text, we cannot know what size the token is. Therefore, we selected samples where 100% tokens in the answer appeared in the OCR text as in Section 6.2.
First, we determine the total area of the tokens in the answer that also appear in the OCR text through the coordinates of the bounding boxes that we used VinVL [60] to extract. We then calculate the percentage of this total area covered in the image. Finally, we determine the group to which these samples belong. With a ratio less than 1%, we classify it into the small group, from 1% to less than 5%, we classify it into the medium group, and the remaining, greater than 5%, we classify it into the large group (see Figure 21).
Figure 22: Results of models based on OCR text size.
As depicted in Figure 22, there is a direct correlation between the increase in OCR text size and the performance of models. As results of the F1-score metric, all models share a common trait that when the text size increases, so does the performance of these models.
Interestingly, although this performance increases as the text is enlarged in the F1-score metric, there is a difference when compared with the EM metric. Specifically, the performance of models, measured by the EM metric, decreases when transitioning from small to medium text sizes but tends to rise slightly again when moving to the large size group.
6.5 Effect of OCR Arrangement
In this research, we conducted an intricate technical analysis to examine the impact of OCR arrangement using three distinct methods on the performance of VQA baseline models applied to the ViTextVQA dataset. Our experimentation involved three sorting approaches: no sorting (maintaining the original order of OCR), sorting based on confidence scores (decrease) extracted from SwintTextSpotter based on the study of Biten et al. [5], and sorting inspired by the top-left to bottom-right strategy as utilized in the PreSTU model by Kil et al. [10].
Table 8: Results of models with different OCR arrangements. $\triangle$ indicates the increase (↑) or decrease (↓) and compares with the original (*).
Model Not Sort Confidence Score Sort Top-Left To Bottom-Right Sort F1-Score (%) EM (%) F1-Score (%) EM (%) F1-Score (%) EM (%) M4C $\triangle$ 25.46* - 10.26* - 18.41 $\downarrow 7.05$ 8.29 $\downarrow 1.97$ 30.04 $\uparrow 4.58$ 11.60 $\uparrow 1.34$ BLIP-2(base) $\triangle$ 36.87* - 13.59* - 35.21 $\downarrow 1.66$ 12.97 $\downarrow 0.62$ 37.78 $\uparrow 0.91$ 15.01 $\uparrow 1.42$ BLIP-2(large) $\triangle$ 38.47* - 15.58* - 36.74 $\downarrow 1.73$ 14.15 $\downarrow 1.43$ 42.16 $\uparrow 3.69$ 18.56 $\uparrow 2.98$ Latr(base) $\triangle$ 40.41* - 17.85* - 38.76 $\downarrow 1.65$ 16.32 $\downarrow 1.53$ 43.13 $\uparrow 2.72$ 20.42 $\uparrow 2.57$ Latr(large) $\triangle$ 42.34* - 20.94* - 39.72 $\downarrow 2.62$ 18.88 $\downarrow 2.06$ 44.07 $\uparrow 1.73$ 22.27 $\uparrow 1.33$ SaL(base) $\triangle$ 43.49* - 20.11* - 43.66 $\uparrow 0.17$ 20.28 $\uparrow 0.17$ 44.89 $\uparrow 1.40$ 20.97 $\uparrow 0.86$ SaL(large) $\triangle$ 43.39* - 21.48* - 43.71 $\uparrow 0.32$ 20.47 $\downarrow 1.01$ 44.74 $\uparrow 1.35$ 21.18 $\downarrow 0.30$ PreSTU(base) $\triangle$ 41.74 $\downarrow 2.07$ 19.15 $\downarrow 1.70$ 40.09 $\downarrow 3.72$ 17.74 $\downarrow 3.11$ 43.81* - 20.85* - PreSTU(large) $\triangle$ 42.85 $\downarrow 2.08$ 21.40 $\downarrow 1.24$ 40.13 $\downarrow 4.80$ 18.43 $\downarrow 4.21$ 44.93* - 22.64* - Avg 39.46 17.81 37.38 16.39 41.73 19.28
Experimental results have shown different performance measures between VQA models when applying different alignment methods for OCR. For details, see Table 8, numbers marked with an asterisk (*) are understood to be the results of the original model.
Top-left to bottom-right sort: All models exhibit an upward performance trend when employing the top-left to bottom-right sorting method. This phenomenon can be attributed to the linguistic characteristics of the Vietnamese language, where characters are typically organized from left to right and top to bottom and accurately reflect the way Vietnamese people read and understand text. This sequential arrangement seems to enhance the semantic context of OCR text, helping the models can easily understand and process information from text in the most natural way.
Confidence score sort: Results found that sorting OCR according to confidence score reduces performance compared to not sort and top-left to bottom-right sort. It indicates that using this strategy may disturb the position of tokens in OCR text, cause the semantic context of OCR text to be lost, and create difficulties for the VQA model in reading and understanding the content.
SaL model: This model exhibits a relatively stable performance regardless of whether it is sorted by top-left to bottom-right or confidence score or not sort. This phenomenon can be attributed to the distinctive characteristics of its two specialized modules, TSS and SCP. These modules provide the model with the capability to pay attention to tokens that have actual significance within OCR text. Consequently, it becomes evident that the SaL model is not heavily reliant on the specific location of tokens in OCR text. This characteristic reduces the dependence of the model on the order of OCR, contributing to its consistent performance across different arrangement types.
In general, arranging OCR text from top-left to bottom-right appears to be a beneficial approach for VQA models when dealing with Vietnamese text. On the contrary, sorting based on confidence scores can introduce ambiguity in placement and result in decreased performance. However, the SaL model stands out as an exception due to the unique capabilities of its two special modules, allowing it to focus on crucial text irrespective of its specific location.
7 Discussion on Characteristics Language for Vietnamese Text-based Visual Question Answering
For a language like Vietnamese, there are unique characteristics that need to be considered to achieve good performance in the VQA task. In this Section, we analyze and discuss two main aspects of the Vietnamese language: Vietnamese diacritics, Vietnamese word segmentation and their relationship to this task.
7.1 Vietnamese Diacritics
The modern Vietnamese language, with its diacritic system, has undergone a complex development process. Initially, ancient Vietnamese did not have diacritics to distinguish phonemes. The influence of Nom and Chinese characters in the history of Vietnamese has contributed to promoting the proposal and development of diacritics. The colonial period marked a great step forward when French intellectuals contributed to the systematization of diacritics. Although there were debates about the necessity of diacritics, they eventually became an indispensable part of Vietnamese, helping to clearly distinguish the sounds and meanings of words.
Today, diacritics are an indispensable part of the Vietnamese national language, including seven letters (ă, â, ê, ô, ơ, ư, đ) and five marks to designate tone(as in à, á, ả, ã, ạ) [67]. They not only distinguish phonemes but also help readers understand and pronounce words correctly. This has made learning and using Vietnamese easier for everyone. An interesting thing is that there are some words in Vietnamese that only differ in tone marks, but have completely different meanings. For example, “má” means “mother”, while “mà” means “but”. This demonstrates the importance of using diacritics in distinguishing words and their meanings.
To underscore the importance of diacritics in Vietnamese, especially in the VQA task, we conducted a series of experiments on our ViTextVQA dataset. In these experiments, we examined diacritic removal levels of 0%, 25%, 50%, 75%, and 100%, respectively, and analyzed the impact of these effects on the performance of the model.
Figure 23: The results of models (base version) based on diacritics remove rate.
The results of the experiments shown in Figure 23 show that removing diacritics significantly affects the ability of models to understand and process language. The performance gradually decreases as the diacritic removal rate increases, as measured by the F1-score and EM. This highlights the role of diacritics in providing important information that helps the models understand questions and predict answers based on the image.
There is a notable difference in the F1-score and EM. In the F1-score, there is observation of a clear and consistent decrease in performance as the diacritic removal rate increases. In contrast, the EM score fluctuates unevenly. Except for the M4C model, the base models experience the steepest decrease at the diacritic removal ratio of 50%, and then have a slight increase or no discernible change as the ratio increases to 75% and 100%.
It is worth noting that the LaTr and SAL models experience a significant drop when the diacritic removal rate increases from 0% to 25%. This suggests that these two models rely heavily on the presence of diacritics.
In general, diacritics not only provide information about context and intonation but also help VQA models understand and interact with the language accurately and effectively.
7.2 Vietnamese Word Segmentation
One exciting aspect of the Vietnamese language is word segmentation, which is complex due to the absence of explicit word boundaries in the written language. Unlike English, white space is a weak indicator of word boundaries in Vietnamese because when written, it is also used to separate syllables that constitute words [68]. The Vietnamese language often forms compound words by combining multiple individual syllables. For example, “thời gian” (“time”) consists of two syllables in a word “thời_gian”. This shows that if the words are segmented exactly, many syllables put together will form a meaningful word.
Figure 24: Distribition of the number of words segmented in OCR text.
We assume that the different arrangement of text from OCR when applying word segmentation will affect the results of the models. That means which arrangement makes more words segmented in OCR text, the more semantic enhancement in context, thereby improving the performance of the models.
To clarify this assumption, we performed word segmentation analysis for different OCR text arrangements. We used the VnCoreNLP tool [59] to perform word segmentation. Refer to Table 9 and Figure 24 for detailed results.
Table 9: Statistic of word segmented in OCR text with different arrangements type.
Arrangement Type Min Word Segmented Max Word Segmented Avg Word Segmented Avg EM (%) Avg F1-Score (%) Not sort 0 11 0.56 17.81 39.46 Confidence score sort 0 10 0.56 16.39 37.38 Top-left to bottom-right sort 0 14 1.06 19.28 41.73
Table 9 clearly shows that top-left to bottom-right sort gives the best results. This is illustrated by the fact that this arrangement achieves the highest average number of segmented words with 1.06 words, the highest number of segmented words in an OCR text at 14 words, and has the best average performance in F1-Score and EM at 41.73 and 19.28 respectively. Arranging OCR text from top-left to bottom-right creates a more systematic context. When words are arranged in order from left to right and top to bottom, the model can easily capture the structure of the text and the relationship between words. This makes word segmentation more accurate, thereby improving the understanding and prediction capabilities of the underlying models.
Focusing on increasing the number of segmented words through OCR text arrangements can be considered a good strategy to improve model performance in many natural language processing applications, especially VQA task, where word-by-word precision plays an important role.
8 Conclusion and Future Work
In this article, we presented a new dataset for the VQA task in Vietnamese to the global research community, especially Vietnam. Specifically, our ViTextVQA dataset includes more than 16,000 images and more than 50,000 question-answer pairs focusing on exploiting text appearing in images. We believe that the effort to create our ViTextVQA dataset will contribute to improving the diversity and enriching resources for the Vietnamese language.
Additionally, we evaluated the performance of state-of-the-art VQA models on our dataset. We found that while these models have proven effective on popular English VQA datasets, similar performance has not been achieved on the ViTextVQA dataset. This prompted us to conduct extensive analyses of the performance aspects of our models, identifying limitations that appear during the experiment, thereby finding valuable insights that lay the groundwork for future research. The results analysis also revealed that reordering OCR text from top-left to bottom-right enhances performance and boosts the accuracy of baseline models in Vietnamese.
In our endeavors moving forward to the future, we recognize the multifaceted utility of our ViTextVQA dataset beyond its application solely in VQA task. One promising avenue lies in leveraging this dataset for the task of Visual Question Generation (VQG). We will aim to use this dataset in VQG task, which the goal is to automatically generate questions about what is happening in the image when given an answer [69, 70, 71].
Additionally, we will continue to explore prompting-based methods with large language models. By combining image features with large language models, prompting-based methods such as [72, 73] have demonstrated a significantly effective, even when trained with fewer parameters than fully fintuning.
Furthermore, our plan also includes developing a chatbot capable of answering image-based questions like Flamingo [1], GPT4 [2], Gemini [3]. This chatbot will be trained on our ViTextVQA dataset and combined with current available VQA datasets in Vietnamese. This will provide an interesting and useful application, helping users interact with the chatbot more naturally and flexibly in Vietnamese.
Acknowledgements
This research is funded by Vietnam National University HoChiMinh City (VNU-HCM) under the grant number DS2024-26-01.
Declarations
Conflict of interest The authors declare that they have no conflict of interest.
Data Availability
Data will be made available on reasonable request.
Appendix
Example Results of Models with OCR Arrangements
Figure 25: Example results of models with not sort.
Figure 26: Example results of models with confidence score sort.
Figure 27: Example results of models with top-left to bottom-right sort.